A gas cylinder holds 0.20 mol of O2 at 170 °C and a pressure of 2.9 atm . The gas expands adiabatically until the volume is doubled. Part A What is the final pressure? Express your answer to two significant figures and include the appropriate units. alue Units Submit Request Answer Part B What is the final temperature in C? Express your answer using two significant figures. oC Submit Request Answer
Answer #1

a) For an adiabatic process

P1V1^\gamma = P2V2^\gamma

\gamma for diatomic gas = 1.4

2.8*V^1.4 = P2*(2V)^1.4

P2 = 1.06 atm

b) For an adiabatic process

T1*V1^(\gamma - 1) = T2*V2^(\gamma - 1)

(273 + 170) * V^(1.4 - 1) = T2*(2V)^(1.4-1)

T2 = 335.73 K

= 335.73 - 273 = 62.73o C